Brenda has 3 more nickles than dimes and 5 fewer pennies than nickles. She has 70 coins in all. How many pennies does she have.: Brenda has 3 more nickles than dimes and 5 fewer pennies than nickles. She has 70 coins in all. How many pennies does she have.
Answer by orca(336)

Let x be the number of dimes.
Then the number of nickels is x + 3,
and the number of pennies is x + 3 - 5 = x - 2.
As the total is 70, we have
x + (x + 3) + (x - 2) = 70
Solving for x, we have
3x + 1 = 70
3x = 69
x = 23
So
the number of dimes is 23,
the number of nickels is x + 3 = 23 + 3 = 26,
the number of pennies is x - 2 = 23 - 2 = 21.